OpenStudy (amistre64):

Is the 'matrix' option in the equation editor just for show? I can never get the thing to operate right and have to type it all in by hand: lets see if this shows :) \[ \begin{array}c \left[ a & b & c \\ d & e & f \\ g & h & i \\ \right] \end{array} \]

OpenStudy (amistre64):

ack!! .... i switched out my \ begin{array}; \ end{array} and my \left, \right

OpenStudy (amistre64):

\[ \left[ \begin{array}c a & b & c \\ d & e & f \\ g & h & i \\ \end{array} \right] \]

OpenStudy (amistre64):

this is what the equation editor does for me: \[(a & b \ c & d)\]

OpenStudy (anonymous):

i get the same thing as amistre64..

OpenStudy (shadowfiend):

Yeah, actually, it is just for show, by which I mean I thought I'd removed it. Actually making matrices work the way they're done in the equation editor turns out to require an actual parser. Most of the `simplification' of LaTeX that happens in the equation editor just consists of allowing you to use commands without the \ . Simpler fractions (e.g., using (5 / 3) for a fraction) and matrices were in the editor a while back, but the parsing turned out to be too slow. I need to revisit that and make it a not-top-down parser, since that requires recursion and it seems like IE's JS is unbelievably bad about recursion. Long story short, yes, I need to disable that until I can revisit the parsing strategy so it doesn't suck, since then I'll be able to actually turn the matrix handling back on.

OpenStudy (amistre64):

var beginning = "\[\left[\begin{array}c " var middle = " ? & ? & ? \\ ? & ? & ? " var ending = "\end{array}\right]\*]" display(middle) function insert() { middle = \text entered\ var matrix = beginning + middle + end } perhaps? :)

OpenStudy (shadowfiend):

Aw hell. Why not. I've been insistent in my mind that we need to make it not suck that much for the beginner, but whatever.